OpenStudy (anonymous):

solve by substitution method. 4x+9y=-5 x=1-3y

OpenStudy (anonymous):

in the top equation, fill in 1-3y for x, which gives \[4\times \left( 1-3y \right)+9y =- 5\]

OpenStudy (anonymous):

gives 4 - 3y = -5, so y =3

OpenStudy (anonymous):

then fill is 3 for y in the lower equation

OpenStudy (anonymous):

which gives \[x = 1-3\times3=-8\]

OpenStudy (anonymous):

it says i need an ordered pair so it would be y=3 and what does x=?

OpenStudy (anonymous):

you take the lower equation: x=1-3y, and substitue y with the number you got from the top equation (y=3)

OpenStudy (anonymous):

ok thanks